OpenStudy (anonymous):

Let f(x) = sqrt(x) a. Find a formula for a function g(x) whose graph is obtained from the graph of f by reflecting across the x-axis, then shifting 2 units up and 4 units to the left. b) Find a formula for a function h(x) whose graph is obtained from f by first shifting 2 units up and 4 units to the right, then reflecting across the x-axis

OpenStudy (anonymous):

I think the first one is g(x) = sqrt(-x+4) + 2

OpenStudy (anonymous):

g(x) = -sqrt(x+4) + 2 ?

OpenStudy (anonymous):

sorry g(x) = -sqrt(x-4) + 2

OpenStudy (anonymous):

If it is on the outside, then it is reflected across the y-axis and on the inside across the x-axis if I remember it right

OpenStudy (anonymous):

you are right in your 1º answer

OpenStudy (anonymous):

can't get it to graph on the calculator with it reflected

OpenStudy (anonymous):

For part b, I have h(x) = -sqrt(x-4) + 2

OpenStudy (anonymous):

first y = sqrt(x) is only valid for x>=0 on reflection with x axis y becomes negative so y = - sqrt(x) then you shift it two units up y = - sqrt(x) + 2 then shift it 4 left so y = - sqrt(x+4) + 2 for second part y = - sqrt(x-4) -2

OpenStudy (anonymous):

so for the first one, is the - inside of the radical and on the second part outside of it?

OpenStudy (luigi0210):

@iambatman

OpenStudy (anonymous):

so my answers for part a and b are g(x) = sqrt(-x+4)+2 h(x) = -(sqrt(x-4))+2

OpenStudy (anonymous):

Look lets start with example if for y = sqrt(x), x=4 then on reflection value of y is inverted, like if initially y=2 for x=4 now y is -2, now if you shift it up then 2 gets added y= - sqrt(x) + 2, shifting left means y = -sqrt(x+4) + 2 eg point (16,4) reflected = (16, -4), displaced upwards (16, -2) displaced left (12, -2) satisfies equation y = -sqrt(x+4) +2 Similarly proceed for the second case

OpenStudy (anonymous):

I remember if you shift left, it is positive and if you shift right it is negative

OpenStudy (anonymous):

what do you mean by displaced upwards and displaced left?